# Cascade adaptive optics with a second stage based on a Zernike wavefront sensor for exoplanet observations

### v1

@@ 2.3 Experimental protocol
In the following, we describe the protocol with the four main steps of our experiments.
@@
The first step consists in injecting phase screens of atmospheric turbulence residuals on the SLM with a frame rate up to 400 Hz. The phase screens have been derived by a numerical simulation using the Object-Oriented Python and Adaptive Optics (OOPAO) simulation tool (Heritier et al. 2023). Here, the first stage AO was correcting for 800 Karhunen–Loève (KL) modes using a PWFS and a 4040 actuator DM using a 6th magnitude star and various wind speed and seeing combinations. The simulated atmospheric turbulence was updated at 2 kHz, but the control loop ran at half that speed by using two subsequent averaged WFS images to update the DM at 1 kHz. The residual phase screens from these simulations at 2 kHz were saved in data cubes to be replayed by the SLM at a speed of 350 Hz. These residuals provide a realistic representation of the atmospheric wavefront errors left after correction by an XAO instrument such as VLT/SPHERE.
@@
The second step deals with the calibration of the ZWFS-based control loop. After setting up the SLM to a flat position, an interaction matrix is first built by sending Hadamard modes (Kasper et al. 2004) to the BMC-492 with 24 actuators across the pupil diameter. The sensor response is recorded on the camera with 36 pixels across the pupil diameter. We then determine the command matrix by calculating the KL modal interaction matrix and computing the pseudo-inverse of this modal interaction matrix for the desired number of modes. This procedure is repeated twice in a row. The first interaction matrix is measured in the presence of aberrations on the Zernike mask. The resulting control matrix is used to close the loop, setting the WFS reference signals to zero. This results in a flat wavefront on the Zernike mask as bench aberrations are compensated by the DM. The corresponding DM position is then used for a second interaction matrix calibration with the ZWFS working around its zero aberration position. This procedure leads to a better calibration of the ZWFS and improved closed loop robustness. We evaluate the effects of both calibration steps in Section A.4.
@@
The third step is related to the temporal control of the AO residuals. The second-stage AO loop is driven with the COSMIC platform (Ferreira et al. 2022), a GPU-based RTC with 2 CPUs and 112 cores in total, and 2 Titan RTX GPUs. It uses a standard AO pipeline, which shows a delay of 110 s between the last pixel received from the WFS and the voltage sent to the DM. The COSMIC Graphics User Interface (GUI) allows us to remotely control the DM, the WFS camera, and the modulation mirror for tests with the PWFS. For the experiments with the ZWFS control loop, no modulation is required and the modulation mirror is left unused in a static position. An additional computer with remote access enables to drive the source illumination, the SLM, the camera in the science arm, and finally the viewer on the wavefront sensing camera with the standalone software. All the functions are remotely accessible independently, and an integrated version of all the software is currently under development.
@@
For the temporal controllers, classical leaky integrator control and reinforcement learning schemes (e.g. Nousiainen et al. 2024) are available with the GHOST interface to explore different control strategies. For the current study with the ZWFS, we limit our temporal control to a classical integrator with a gain and leak to mimic standard AO control with the second-stage AO. Further tests with more elaborated real-time control laws will be considered in the near future. Unless otherwise stated, the ZWFS-based control loop runs at the same speed as the SLM (350 Hz) to simulate a second stage with a speed that is twice the update rate of the first stage.
@@
Finally, the fourth step corresponds to the data acquisition in open and closed loop for the second-stage AO for performance assessment. In the following, we measure the residual wavefront errors from the WFS path to estimate the AO correction. The residual WFS measurement is given in normalized DM commands [V] after multiplication by the control matrix (0-200 V are commanded by values ranging between 0 and 1). We also acquire data on the science arm with or without coronagraph to estimate the contrast difference in the normalized coronagraphic images between the open and closed loop situations. The images are sampled with 3.22 pixel per resolution element.
@@ 3.1 Assumptions
We run a first experiment with the ZWFS-based second-stage AO loop using standard observing conditions to provide a proof of concept. The SLM injects phase screens corresponding to XAO residuals based on VLT/SPHERE characteristics under median observing conditions: a 10 m.s-1 wind speed and 0.7” seeing. The size of the field stop is set to 35 /D.
@@ 3.2 Correction of the AO residuals
After the generation of the command matrix, we operate the ZWFS control loop with a classical integrator for the temporal control law with a gain of 0.8 and a leak of 0.99. In closed loop, the number of corrected KL modes is set to 350. Figure 3 top plot shows the temporal evolution of the wavefront errors for a subset of modes before and after closing the ZWFS control loop, open loop (OL) and closed loop (CL) respectively. A reduction of the temporal standard deviation is observed for all the modes and is mainly noticeable for the low-order modes, in particular the modes 1 and 2 which can be associated to pointing errors.
@@
Figure 4 shows the temporal standard deviation of the wavefront errors for each mode before and after closing the loop in top plot and the effective gain in bottom plot. Our ZWFS-based second stage AO loop reduces the wavefront errors by at least a factor of 5 for all the modes and shows a gain up to about 13 for the low-order modes. In closed loop, the wavefront errors slightly increase with the highest modes and studies are on going to understand this point further. The observed effect could find its origin in the small difference in signal sampling between the ZWFS and the DM, respectively 36 pixels and 24 actuators across the pupil diameter. In this configuration, the transfer function of the sensor pixel introduces some sensitivity loss in the control loop for the modes with large spatial frequency content and additional calibration may be required to address this undesired feature. Still, this effect remains negligible in comparison with the large wavefront error reduction between the open and closed loop operations.
@@
The wavefront error reduction is also visible when we consider the quadratic sum of all the modes to analyze the total amount of aberrations in open and closed loop, see Fig. 3 bottom plot. We estimate the temporal mean of the RMS wavefront error and of (7.40.6) and (1.20.3)V in open and closed loop, showing an overall reduction of the wavefront errors by a factor of 6 and a decrease in the temporal wavefront error dispersion by a factor of 2, confirming the ability of the ZWFS control loop to reduce the AO residuals and increase the wavefront error stability. In closed loop, we also observe a 1 s cyclic behavior of the total wavefront errors which is also clearly present in Mode 2. This 1 Hz appears to be a beat frequency since zooming in, there seems to be a very high frequency pattern, close to the Nyquist frequency. Actually, our analysis shows a good fit between the Mode 2 data and a sine wave of period 1.6 ms (625 Hz), multiplied by a squared cosine of period 1.6 seconds, producing the beating period of 0.8s (1.2 Hz). This would correspond to the sum of two sine waves separated in frequency by 2.4 Hz. We are currently investigating the origin of this effect, which could be related to vibrations in the bench or undesired effects in the temporal control.
@@
We also explore the temporal behavior of our second-stage AO loop by representing the temporal power spectral density (PSD) of the residual aberrations in open and closed loops, see Fig. 5. Our loop enables to reduce aberrations by at least two orders of magnitude for wavefront errors with temporal frequencies up to 10 Hz and larger than 1 for phase aberrations with temporal frequencies up to about 30 Hz. The 1 Hz peak observed in closed loop corresponds to the cyclic behavior that is reported in Fig. 3 bottom plot. At larger temporal frequencies, the PSD shows slightly higher residuals, underlining some artifacts introduced by our control loop. However, the wavefront errors are small enough at these temporal frequencies to have no impact on the overall wavefront correction.
@@ 3.3 Contrast improvement in coronagraphic images
We now study the impact of the AO residuals correction into contrast improvement in the coronagraphic image of the source. Our experiment is performed with a CLC and in the presence of NCPA between the sensing path and the science arm which are left uncorrected in this experiment, thus limiting the ultimate achievable contrast with our control loop.
Figure 6 top plot shows the recorded coronagraphic images in the science camera in open and closed loops. The anatomy of such images has been described in the literature (e.g. Guyon 2018; Cantalloube et al. 2019) and here, we briefly recall the origin of the main features. The images exhibit a large high-contrast region which radius corresponds to the spatial frequency cutoff of the first stage XAO wavefront correction. Located at the edge of the images, some bright speckles outside the control region are produced by the DM actuator print-through. Compared to the open loop image, the closed-loop image shows the formation of an additional high-contrast region inside the XAO control region, at a distance from the source corresponding to a spatial frequency of about 11 cycle/pupil (c/pup). This value is consistent with the expected frequency cutoff of the second-stage AO loop with the ZWFS for the correction of 350 KL modes (), confirming the effective correction performed by our control loop.
Figure 6 middle and bottom plots show the azimuthally averaged intensity profiles of the coronagraphic images in open and closed loops and the estimated contrast gain between them. We observe a contrast improvement with a gain by a factor up to 2 in the ZWFS-based controlled region ranging from 2 to 11 , showing the ability of our control loop to improve the contrast in the coronagraphic image. In the middle plot, we also represent the azimuthally averaged intensity profile of the coronagraphic in the absence of XAO residuals on the SLM, representing the ultimate contrast floor in our experiment. Our control loop manages to achieve this threshold at the separations out to 8 . This contrast floor in our images finds several explanations, the main limitations in our experiment are most likely the limited contrast provided by the CLC and the presence of the NCPA between the wavefront sensing path and the science arm which lead to quasi-static speckles that alter the image contrast. Based on the measurements with the Fast & Furious phase diversity algorithm (Keller et al. 2012; Korkiakoski et al. 2014; Wilby et al. 2018; Bos et al. 2020) on the science camera, we estimate a total amount of NCPA of 20 nm RMS for the first 36 Zernike modes after piston.
@@ 4.1 Wind speed
In XAO instruments, the observation of planetary companions or disks around an imaged star can be altered by the presence of a bright stellar veil called wind-driven halo (Cantalloube et al. 2018, 2020). This harmful feature is related to bad observing conditions for which the AO loop runs slower than the atmospheric turbulence evolution. One of the main drivers is the wind speed which varies in amplitude and direction with altitude and time. Under these observing conditions, a single-stage XAO sees its atmospheric aberration compensation limited by the large temporal error between the wavefront error measurement and correction in the AO control loop. The so-called AO servo-lag error can be reduced at (i) the software level, e.g. by using predictive control strategies or PSF reconstruction for post-processing or at (ii) the hardware level with the use of a faster AO control loop. For the latter, a second-stage AO with a faster control loop constitutes an attractive solution to reduce the temporal error and address the residuals left from the first XAO stage. This approach is here assessed with our ZWFS-based control loop.
@@
We repeat a similar experiment with different (5 and 10 m.s-1) and of 1.0” and observe the same trends in terms of contrast gain and threshold level, although at a natural less deep contrast. At a larger wind speed (=34 m.s-1), the control loop breaks down and no more correction is applied. The first stage AO residuals are too large for the ZWFS to remain in the linear range. Sign changes in the reconstructed wavefront errors will result in rapid loop divergence. The stability of the control loop is recovered by reducing down to 250 modes.
@@ 4.2 Seeing
Together with wind speed, another key feature of the atmospheric turbulence conditions is the seeing. This term denotes the PSF full width at half maximum, which increases with the spatio-temporal fluctuations of the atmospheric perturbation, leading to large wavefront errors to compensate for. At large seeings, these wavefront errors prove too large to be fully corrected for by a single XAO, limiting the contrast in the coronagraphic images of an observed star. To overcome the contrast threshold due to the atmospheric turbulence, addressing the atmospheric residuals with a second-stage AO is again a promising solution to enlarge the science return of high-contrast facilities over a wider range of seeing conditions. We here investigate the capabilities of our ZWFS-based control loop for different seeing conditions.
@@
Our scheme achieves the largest contrast gain for the largest values of corresponding to conditions in which the wavefront error residuals from the first AO loop are the more important. For instance, a gain by a factor of up to 3.3 is observed at a separation of 3 for of 1.0”. These results emphasize the benefit of having a second-stage AO loop with a faster correction to compensate for the XAO residuals when the observing conditions are less favorable and thus, expand the science return for exoplanet imaging. The ZWFS-based control loop still offers a contrast gain up to 2 at 3 for the best seeing with =0.5”, enabling access to the faintest planets in excellent observing conditions.
In our experiments, some instabilities in the control loop are observed for of 1.0”, leading to a loop break. These effects are removed by simply reducing from 350 to 300. Our ZWFS control loop runs efficiently in different seeings by simply adjusting the number of corrected KL modes.
@@ 4.3 Source flux
We study the impact of the source flux on the wavefront error correction with our control loop. The source flux is adjusted with the current applied on the SLED source control. Our previous experiments are conducted in high flux regime for which is set to 25 A, serving as a reference in terms of source flux. In the following, we estimate the integrated flux in the core of the non coronagraphic image within a photometric aperture of 2.44 /D diameter for a given . The source flux is then expressed with respect to the reference source flux in relative magnitude mag.
Figure 9 shows the intensity profiles of the coronagraphic images in open and closed loops for different mag. The contrast performance is stable for mag up to 2.9. At larger relative magnitudes, the contrast degrades in the second-stage controlled region, showing a loss of efficiency of our scheme. In control loop operation, this effect relates to noise amplification at all the spatial frequencies, translating into contrast degradation in the area of the coronagraphic image within the ZWFS-based control radius. The result allows for an estimate of the source flux limit at which our second-stage AO scheme operates efficiently.
To address the noise amplification at the low flux regime, we perform tests on our control loop by adjusting the integrator gain. Such a procedure replicates the on-sky operation with a real AO system to ensure a stable closed loop in low flux regime. Figure 10 shows an example for different loop gains with mag of 5.04. The noise amplification in the coronagraphic images for an integrator gain of 0.80 is removed by adjusting the loop gain to 0.15, leading a contrast gain that is comparable to the one obtained at the highest source flux in our experiments (mag=0.0). Our ZWFS control loop works in low flux regime and possible noise amplification is eliminated by adjusting the integrator gain.

### v2

@@ 2.3 Experimental protocol
In the following, we describe the protocol with the four main steps of our experiments. The first step consists of injecting phase screens of atmospheric turbulence residuals on the SLM with a frame rate up to 400 Hz. The phase screens have been derived by a numerical simulation using the Object-Oriented Python and Adaptive Optics (OOPAO) simulation tool (Heritier et al. 2023). Here, the first stage AO was correcting for 800 Karhunen–Loève (KL) modes using a PWFS and a 4040 actuator DM using a sixth-magnitude star and various wind speed and seeing combinations. The simulated atmospheric turbulence was updated at 2 kHz, but the control loop ran at half that speed by using two subsequent averaged WFS images to update the DM at 1 kHz. The residual phase screens from these simulations at 2 kHz were saved in data cubes to be replayed by the SLM at a speed of 350 Hz. These residuals provide a realistic representation of the atmospheric wavefront errors left after correction by an XAO instrument such as VLT/SPHERE.
@@
The second step deals with the calibration of the ZWFS-based control loop. After setting up the SLM to a flat position, an interaction matrix is first built by sending Hadamard modes (Kasper et al. 2004) to the BMC-492 with 24 actuators across the pupil diameter. The sensor response is recorded on the camera with 36 pixels across the pupil diameter. We then determine the command matrix by calculating the KL modal interaction matrix and computing the pseudo-inverse of this modal interaction matrix for the desired number of modes. This procedure is repeated twice in a row. The first interaction matrix is measured in the presence of aberrations on the Zernike mask. The resulting control matrix is used to close the loop, setting the WFS reference signals to zero. This results in a flat wavefront on the Zernike mask as bench aberrations are compensated by the DM. The corresponding DM position is then used for a second interaction matrix calibration with the ZWFS working around its zero-aberration position. This procedure leads to a better calibration of the ZWFS and improved closed loop robustness. We evaluate the effects of both calibration steps in Sect. A.4.
@@
The third step is related to the temporal control of the AO residuals. The second-stage AO loop is driven with the COSMIC platform (Ferreira et al. 2022), a GPU-based RTC with 2 CPUs and 112 cores in total, and 2 Titan RTX GPUs. It uses a standard AO pipeline, which shows a delay of 110 s between the last pixel received from the WFS and the voltage sent to the DM. The COSMIC Graphics User Interface (GUI) allows us to remotely control the DM, the WFS camera, and the modulation mirror for tests with the PWFS. For the experiments with the ZWFS control loop, no modulation is required and the modulation mirror is left unused in a static position. An additional computer with remote access allows us to drive the source illumination, the SLM, the camera in the science arm, and finally the viewer on the wavefront sensing camera with the standalone software. All the functions are remotely accessible independently, and an integrated version of all the software is currently under development.
@@
For the temporal controllers, classical leaky integrator control and reinforcement learning schemes (e.g., Nousiainen et al. 2024) are available with the GHOST interface to explore different control strategies. For the current study with the ZWFS, we limit our temporal control to a classical integrator with a gain and leak to mimic standard AO control with the second-stage AO. Further tests with more elaborate real-time control laws will be considered in the near future. Unless otherwise stated, the ZWFS-based control loop runs at the same speed as the SLM (350 Hz) to simulate a second stage with a speed that is twice the update rate of the first stage.
@@
Finally, the fourth step corresponds to the data acquisition in open and closed loop for the second-stage AO for performance assessment. In the following, we present our measurements of the residual wavefront errors from the WFS path to estimate the AO correction. The residual WFS measurement is given in normalized DM commands [V] after multiplication by the control matrix (0-200 V are commanded by values ranging between 0 and 1). We also acquired data on the science arm with or without coronagraph to estimate the contrast difference in the normalized coronagraphic images between the open and closed loop situations. The images were sampled with 3.22 pixel per resolution element.
@@ 3.1 Assumptions
We ran a first experiment with the ZWFS-based second-stage AO loop using standard observing conditions to provide a proof of concept. The SLM injects phase screens corresponding to XAO residuals based on VLT/SPHERE characteristics under median observing conditions: a 10 m.s-1 wind speed and 0.7” seeing. The size of the field stop was set to 35 /D.
@@ 3.2 Correction of the AO residuals
After the generation of the command matrix, we operated the ZWFS control loop with a classical integrator for the temporal control law with a gain of 0.8 and a leak of 0.99. In closed loop, the number of corrected KL modes is set to 350. Figure 3 top plot shows the temporal evolution of the wavefront errors for a subset of modes before and after closing the ZWFS control loop, open loop (OL) and closed loop (CL) respectively. A reduction of the temporal standard deviation is observed for all the modes and is mainly noticeable for the low-order modes, in particular the modes 1 and 2, which can be associated with pointing errors.
@@
Figure 4 shows the temporal standard deviation of the wavefront errors for each mode before and after closing the loop in top plot and the effective gain in the bottom plot. Our ZWFS-based second stage AO loop reduces the wavefront errors by at least a factor of 5 for all the modes and shows a gain up to about 13 for the low-order modes. In closed loop, the wavefront errors slightly increase with the highest modes and studies are on going to understand this point further. The observed effect could find its origin in the small difference in signal sampling between the ZWFS and the DM, respectively 36 pixels and 24 actuators across the pupil diameter. In this configuration, the transfer function of the sensor pixel introduces some sensitivity loss in the control loop for the modes with large spatial frequency content and additional calibration may be required to address this undesired feature. Still, this effect remains negligible in comparison with the large wavefront error reduction between the open and closed loop operations.
@@
The wavefront error reduction is also visible when we consider the quadratic sum of all the modes to analyze the total amount of aberrations in open and closed loop (see Fig. 3, bottom plot). We estimate the temporal mean of the RMS wavefront error and of (7.40.6) and (1.20.3)V in open and closed loop, showing an overall reduction of the wavefront errors by a factor of 6 and a decrease in the temporal wavefront error dispersion by a factor of 2, confirming the ability of the ZWFS control loop to reduce the AO residuals and increase the wavefront error stability. In closed loop, we also observe a 1 s cyclic behavior of the total wavefront errors which is also clearly present in Mode 2. This 1 Hz appears to be a beat frequency since zooming in, there seems to be a very high frequency pattern, close to the Nyquist frequency. Actually, our analysis shows a good fit between the Mode 2 data and a sine wave of period 1.6 ms (625 Hz), multiplied by a squared cosine of period 1.6 seconds, producing the beating period of 0.8s (1.2 Hz). This would correspond to the sum of two sine waves separated in frequency by 2.4 Hz. We are currently investigating the origin of this effect, which could be related to vibrations in the bench or undesired effects in the temporal control.
@@
We also explored the temporal behavior of our second-stage AO loop by representing the temporal power spectral density (PSD) of the residual aberrations in open and closed loops, see Fig. 5. Our loop enabled us to reduce aberrations by at least two orders of magnitude for wavefront errors with temporal frequencies up to 10 Hz and larger than 1 for phase aberrations with temporal frequencies up to about 30 Hz. The 1 Hz peak observed in closed loop corresponds to the cyclic behavior that is reported in Fig. 3 (bottom plot). At higher temporal frequencies, the PSD shows slightly higher residuals, underlining some artifacts introduced by our control loop. However, the wavefront errors are small enough at these temporal frequencies to have no impact on the overall wavefront correction.
@@ 3.3 Contrast improvement in coronagraphic images
We go on to study the impact of the AO residuals correction into contrast improvement in the coronagraphic image of the source. Our experiment was performed with a CLC and in the presence of NCPA between the sensing path and the science arm, which are left uncorrected in this experiment; thus, this limits the ultimate achievable contrast with our control loop.
Figure 6 top plot shows the recorded coronagraphic images in the science camera in open and closed loops. The anatomy of such images has been described in the literature (e.g., Guyon 2018; Cantalloube et al. 2019) and here, we briefly recall the origin of the main features. The images exhibit a large high-contrast region whose radius corresponds to the spatial frequency cutoff of the first stage XAO wavefront correction. Located at the edge of the images, some bright speckles outside the control region are produced by the DM actuator print-through. Compared to the open loop image, the closed-loop image shows the formation of an additional high-contrast region inside the XAO control region, at a distance from the source corresponding to a spatial frequency of about 11 cycle/pupil (c/pup). This value is consistent with the expected frequency cutoff of the second-stage AO loop with the ZWFS for the correction of 350 KL modes (), confirming the effective correction performed by our control loop.
Figure 6 (middle and bottom plots) shows the azimuthally averaged intensity profiles of the coronagraphic images in open and closed loops and the estimated contrast gain between them. We observe a contrast improvement with a gain by a factor of up to 2 in the ZWFS-based controlled region ranging from 2 to 11 , showing the ability of our control loop to improve the contrast in the coronagraphic image. In the middle plot, we also represent the azimuthally averaged intensity profile of the coronagraphic in the absence of XAO residuals on the SLM, representing the ultimate contrast floor in our experiment. Our control loop manages to achieve this threshold at the separations out to 8 . This contrast floor in our images finds several explanations, the main limitations in our experiment are most likely the limited contrast provided by the CLC and the presence of the NCPA between the wavefront sensing path and the science arm which lead to quasi-static speckles that alter the image contrast. Based on the measurements with the Fast & Furious phase diversity algorithm (Keller et al. 2012; Korkiakoski et al. 2014; Wilby et al. 2018; Bos et al. 2020) on the science camera, we estimate a total amount of NCPA of 20 nm RMS for the first 36 Zernike modes after piston.
@@ 4.1 Wind speed
In XAO instruments, the observation of planetary companions or disks around an imaged star can be altered by the presence of a bright stellar veil called wind-driven halo (Cantalloube et al. 2018, 2020). This harmful feature is related to bad observing conditions for which the AO loop runs slower than the atmospheric turbulence evolution. One of the main drivers is the wind speed which varies in amplitude and direction with altitude and time. Under these observing conditions, a single-stage XAO sees its atmospheric aberration compensation limited by the large temporal error between the wavefront error measurement and correction in the AO control loop. The so-called AO servo-lag error can be reduced at (i) the software level, for instance, by using predictive control strategies or PSF reconstruction for post-processing or at (ii) the hardware level with the use of a faster AO control loop. For the latter, a second-stage AO with a faster control loop constitutes an attractive solution to reduce the temporal error and address the residuals left from the first XAO stage. This approach is here assessed with our ZWFS-based control loop.
@@
We repeated a similar experiment with different (5 and 10 m.s-1) and of 1.0” and observed the same trends in terms of contrast gain and threshold level, although at a natural less deep contrast. At a higher wind speed (=34 m.s-1), the control loop breaks down and no more correction is applied. The first stage AO residuals are too large for the ZWFS to remain in the linear range. Sign changes in the reconstructed wavefront errors will result in rapid loop divergence. The stability of the control loop is recovered by reducing down to 250 modes.
@@ 4.2 Seeing
Together with wind speed, another key feature of the atmospheric turbulence conditions is the seeing. This term denotes the PSF full width at half maximum, which increases with the spatio-temporal fluctuations of the atmospheric perturbation, leading to large wavefront errors to compensate for. At large seeings, these wavefront errors prove too large to be fully corrected for by a single XAO, limiting the contrast in the coronagraphic images of an observed star. To overcome the contrast threshold due to the atmospheric turbulence, addressing the atmospheric residuals with a second-stage AO is again a promising solution to enlarge the science return of high-contrast facilities over a wider range of seeing conditions. Here, we investigate the capabilities of our ZWFS-based control loop for different seeing conditions.
@@
Our scheme achieves the largest contrast gain for the largest values of corresponding to conditions in which the wavefront error residuals from the first AO loop are the more important. For instance, a gain by a factor of up to 3.3 is observed at a separation of 3 for of 1.0”. These results emphasize the benefit of having a second-stage AO loop with a faster correction to compensate for the XAO residuals when the observing conditions are less favorable, thereby extending the science return for exoplanet imaging. The ZWFS-based control loop still offers a contrast gain up to 2 at 3 for the best seeing with =0.5”, enabling access to the faintest planets in excellent observing conditions.
In our experiments, some instabilities in the control loop are observed for of 1.0”, leading to a loop break. These effects can be removed by simply reducing from 350 to 300. Our ZWFS control loop runs efficiently in different seeings by simply adjusting the number of corrected KL modes.
@@ 4.3 Source flux
We studied the impact of the source flux on the wavefront error correction with our control loop. The source flux was adjusted with the current applied on the SLED source control. Our previous experiments were conducted in high flux regime for which is set to 25 A, serving as a reference in terms of source flux. In the following, we present our estimates of the integrated flux in the core of the non-coronagraphic image within a photometric aperture of 2.44 /D diameter for a given . The source flux can then be expressed with respect to the reference source flux in relative magnitude, mag.
Figure 9 shows the intensity profiles of the coronagraphic images in open and closed loops for different mag. The contrast performance is stable for mag up to 2.9. At larger relative magnitudes, the contrast degrades in the second-stage controlled region, showing a loss of efficiency for our scheme. In control loop operation, this effect relates to noise amplification at all the spatial frequencies, translating into contrast degradation in the area of the coronagraphic image within the ZWFS-based control radius. The result allows for an estimate of the source flux limit at which our second-stage AO scheme can operate efficiently.
To address the noise amplification at the low flux regime, we performed tests on our control loop by adjusting the integrator gain. Such a procedure replicates the on-sky operation with a real AO system to ensure a stable closed loop in low flux regime. Figure 10 shows an example for different loop gains with mag of 5.04. The noise amplification in the coronagraphic images for an integrator gain of 0.80 is removed by adjusting the loop gain to 0.15, leading a contrast gain that is comparable to the one obtained at the highest source flux in our experiments (mag=0.0). Our ZWFS control loop works in low flux regime and possible noise amplification is eliminated by adjusting the integrator gain.